# Loss Gradient Gaussian Width based Generalization and Optimization Guarantees

### v1

@@ Abstract
Generalization and optimization guarantees on the population loss in machine learning often rely on uniform convergence based analysis, typically based on the Rademacher complexity of the predictors. The rich representation power of modern models has led to concerns about this approach. In this paper, we present generalization and optimization guarantees in terms of the complexity of the gradients, as measured by the Loss Gradient Gaussian Width (LGGW). First, we introduce generalization guarantees directly in terms of the LGGW under a flexible gradient domination condition, which we demonstrate to hold empirically for deep models. Second, we show that sample reuse in finite sum (stochastic) optimization does not make the empirical gradient deviate from the population gradient as long as the LGGW is small. Third, focusing on deep networks, we present results showing how to bound their LGGW under mild assumptions. In particular, we show that their LGGW can be bounded (a) by the $L_{2}$-norm of the loss Hessian eigenvalues, which has been empirically shown to be $\tilde{O}(1)$ for commonly used deep models; and (b) in terms of the Gaussian width of the featurizer, i.e., the output of the last-but-one layer. To our knowledge, our generalization and optimization guarantees in terms of LGGW are the first results of its kind, avoid the pitfalls of predictor Rademacher complexity based analysis, and hold considerable promise towards quantitatively tight bounds for deep models.
@@ 1 Introduction
Machine learning theory typically characterizes generalization behavior and related aspects in terms of uniform convergence guarantees, often in terms of Rademacher complexities of predictor function classes [44, 9, 73, 60]. However, given the rich representation power of modern deep models, it has been difficult to develop sharp uniform bounds, and concerns have been raised whether non-vacuous bounds following the uniform convergence route may be possible [61, 62, 43, 42, 11, 31]. On the other hand, the empirical successes of deep learning has highlighted the importance of gradients, which play a critical role in learning over-parameterized models with deep representations [14, 32, 45]. The motivation behind this paper is to investigate generalization and optimization guarantees based on uniform convergence analysis of gradients, rather than the predictors.
@@
Empirical work in recent years has demonstrated that gradients of overparameterized deep learning models are typically “simple,” e.g., only a small fraction of entries have large values [64, 65, 47, 86, 88], gradients typically lie in a low-dimensional space [33, 29, 71, 39], using top-$k$ components with suitable bias corrections work well in practice [71, 39], etc. In this paper, we view such simplicity in terms of the loss gradient Gaussian width (LGGW) [80, 83]. Based on the empirical evidence, we posit that loss gradients of modern machine learning models have small Gaussian widths. Our main technical results in the paper present generalization and optimization guarantees for learning in terms of the LGGW. While our main technical results hold irrespective of whether the LGGW is small or not, similar to their classical Rademacher complexity counterparts, the results imply sharper bounds if the LGGW is small. Further, while deep learning gradients form the motivation behind the work, some of the technical results we present are not restricted to deep learning and hold for general models under simple regularity conditions.
@@
In Section 2, we present generalization bounds in terms of the LGGW. Our analysis needs a mechanism to transition from losses to gradients of losses, and we assume the population loss function satisfies a flexible gradient domination (GD) condition [27], which includes the PL (Polyak-Łojasiewicz) condition [66] used in modern deep learning [50, 53, 49] as a simple special case. We empirically demonstrate that GD holds for commonly used deep models. Then, based on recent advances in uniform convergence of gradients and, more generally, vector-valued functions [56, 57, 27], one can readily get generalization bounds based on the vector Rademacher complexity of gradients. Our main new result on generalization is to show that such vector Rademacher complexity can be bounded by the LGGW, which makes the bound explicitly depend on the geometry of gradients. The result may seem surprising at first as the expectation of randomness in vector Rademacher complexity is over $n$ Rademacher variables $\{\varepsilon_{i},i\in[n]\}$ corresponding to the samples whereas the expectation of randomness in Gaussian width is over $p$ Gaussian variables $\mathbf{g}\sim{\cal N}(0,\mathbb{I}_{p\times p})$ corresponding to the $p$ parameters in the model. Our proof is based on generic chaining (GC) [80], where we first bound the vector Rademacher complexity using a hierarchical covering argument, and then bound the hierarchical covering in terms of the Gaussian width using the majorizing measure theorem [80]. To our knowledge, this is the first generalization bound in terms of Gaussian width of loss gradients, and the results imply that models with small LGGW will generalize well.
@@
In Section 3, we present optimization guarantees based on the LGGW. One of the arguably biggest disconnects between the theory and practice of (mini-batch) stochastic gradient descent (SGD) is that the theory assumes a mini-batch of fresh samples in each step whereas in practice one reuses samples. If one just focuses on the finite sum empirical risk minimization (ERM), sample reuse is not an issue. However, the goal of such optimization is to reduce the population loss, and sample reuse derails the standard analysis. We show that even with sample reuse, the discrepancy between the (mini-batch) sample average and the population gradient can be bounded by the LGGW along with only a logarithmic dependence on how many times the sample has been reused. Thus, if the LGGW is small, the sample average gradients stay close to the population gradients, even under sample reuse. We establish such results for both full-batch and mini-batch, and use these results to establish population convergence results for mini-batch gradient descent with sample reuse. Since our bounds are uniform, there is no dependence on the mechanism by which the mini-batches are selected, e.g., random draw with replacement [34, 67], permute/shuffle and draw without replacement [54, 72], etc.
In Section 4, we present the first results on bounding LGGW of deep learning gradients. Recent empirical work has demonstrated sharply decaying eigenvalues of the loss Hessian [92, 87, 89]. Based on such observations, our first result is to show that the LGGW can be bounded by an upper bound on the $L_{2}$-norm of the loss Hessian eigenvalues, which has been empirically shown to be $\tilde{O}(1)$ for commonly used deep models  [92, 87]. Further, we consider formal models for feed-forward and residual networks widely studied in the deep learning theory literature [1, 21, 7, 51, 8] and show that their LGGW can be bounded by the Gaussian width of the featurizer, i.e., the output of the last-but-one layer. The results emphasize the benefits of Hessian eigendecay and the simplicity as well as stability of featurizers for deep learning.
@@
A brief word on notation: $c_{1},c_{2},c_{3}$, etc., denote constants, and they may mean different constants at different places. Such repeat use of constants is common in the GC literature [80]. We use $\tilde{O}$ to hide poly-log terms. The exact dependencies are in the proofs in the Appendix.
@@ 2 Gaussian Width based Generalization Bounds
In this section, we formalize the notion of Loss Gradient Gaussian Width (LGGW) [83, 80], and as the main result, establish generalization bounds in terms of LGGW. For parameters $\theta\in\Theta\subset\mathbb{R}^{p}$ and data points $z\in{\cal Z}$ where $z=(\mathbf{x},y),\mathbf{x}\in\mathbb{R}^{d},y\in\mathbb{R}$, for any suitable loss $\ell(\theta;z):\Theta\times{\cal Z}\mapsto\mathbb{R}_{+}$, the gradients of interest are $\xi(\theta;z)=\nabla_{\theta}\ell(\theta;z)\in\mathbb{R}^{p}$. Given a dataset of $n$ samples, we start by defining the sets of all empirical average gradients and $n$-tuple of individual gradients:
@@
Given $n$ samples $z^{(n)}=\{z_{1},\ldots,z_{n}\}\in{\cal Z}^{(n)}$, for a suitable choice of parameter set $\Theta$ and loss $\ell:\Theta\times{\cal Z}\mapsto\mathbb{R}_{+}$, let the set of all possible empirical average and $n$-tuple of individual loss gradients be respectively denoted as:
$\hat{\Xi}_{n}(\Theta)$ :=$:=\left\{\hat{\xi}_{n}\in\mathbb{R}^{p}\mid\exists\theta\in\Theta~{}~{}~{}\text{s.t.}~{}~{}\hat{\xi}_{n}=\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta;z_{i})\right\}~{},$ (1)
$\hat{\Xi}^{(n)}(\Theta)$ :=$:=\left\{\hat{\xi}^{(n)}=\{\xi_{i},i\in[n]\}\mid\exists\theta\in\Theta~{}\text{s.t.}~{}\xi_{i}=\nabla\ell(\theta;z_{i})\right\}~{}.$ (2)
For convenience, we will denote the empirical average gradient in (1) as $\hat{\Xi}_{n}$, and the set of $n$-tuples of gradients in (2) as $\hat{\Xi}^{(n)}$. Note that the sets $\hat{\Xi}_{n}$ and $\hat{\Xi}^{(n)}$ are conditioned on $z^{(n)}$, and the results we will present will hold for any $z^{(n)}\subset{\cal Z}^{n}$, i.e., any $n$ samples under suitable regularity conditions (see Section 4.2.1). We use $\hat{\xi}_{n}(\theta)$ and $\hat{\xi}_{n}$ interchangeably to denote the n sample average gradient. Next, we define the Loss Gradient Gaussian Width:
@@
Given the empirical average gradient set $\hat{\Xi}_{n}$ as in Definition 1, the loss gradient Gaussian width (LGGW) is defined as:
$w(\hat{\Xi}_{n}):=\mathbb{E}_{\mathbf{g}\sim{\cal N}(0,\mathbb{I}_{p\times p})}\left[\sup_{\hat{\xi}_{n}\in\hat{\Xi}_{n}}~{}\langle\hat{\xi}_{n},\mathbf{g}\rangle\right]~{}.$ (3)
@@ 2.1 Bounds based on Vector Rademacher Complexity
Our approach to generalization bounds is based on vector Rademacher complexities. Unlike the traditional approach which focus on bounds based on Rademacher complexities of the set of predictors $f(\theta;x)$ [44, 9, 73, 60], we will develop generalization bounds based on vector Rademacher complexities [56, 57, 27] of the set of loss gradients $\xi(\theta,z)=\nabla\ell(\theta;z)$. In particular, following [27], we consider Normed Empirical Rademacher Complexity (NERC), which is the extension of standard Rademacher complexity to sets of vectors or vector-valued functions.
@@
$\min_{\theta}{\cal L}_{{\cal D}}(\theta):=\mathbb{E}_{z\sim{\cal D}}\left[\ell(\theta;z)\right].$ (5)
The learner does not observe the distribution ${\cal D}$ and typically minimize the empirical risk
^$\hat{\cal L}_{n}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\ell(\theta;z_{i}),$ (6)
@@
It is practically important is to understand whether the $(\alpha,\bar{c}_{\alpha})$-GD condition in Assumption 1 holds for widely used models, e.g., modern deep learning models. In Figure 1, we plot the gradient domination ratio $\frac{{\cal L}_{\mathcal{D}}(\theta)}{\left\|\nabla{\cal L}_{\mathcal{D}}(\theta)\right\|^{\alpha}}$ for $\alpha=1,2$ for three standard deep learning models. The plots demonstrate that the condition indeed clearly holds for $\alpha=1$ with a small constant $\bar{c}_{1}<5$ for all models, and for $\alpha=2$, i.e., the PL condition [66, 50, 53, 49], with $\bar{c}_{2}<1$ for some models, but needs large constants $\bar{c}_{2}>100$ for for ResNet18 on CIFAR-10. We establish results for any $\alpha\in[1,2]$, which along with the empirical evidence for $\alpha=1$ makes the Assumption 1 valid. ∎
@@
Figure 1: Gradient Domination Ratio for $\alpha=1,2$ for (a) ResNet18 on CIFAR-10, (b) FFN on CIAFR-10 and (c) CNN on Fashion-MNIST. Each experiment is repeated 5 times and we report the average and the maximum ratio. The results demonstrate that the gradient domination condition indeed clearly holds for $\alpha=1$ with a small constant $\bar{c}_{1}<5$ for all models, and for $\alpha=2$, with $\bar{c}_{2}<1$ for some models, but needs large constants $\bar{c}_{2}>100$ for for ResNet18.
@@
${\cal L}_{{\cal D}}(\hat{\theta})-{\cal L}_{{\cal D}}(\theta^{*})\leq 2\bar{c}_{\alpha}\bigg{(}\|\nabla\hat{{\cal L}}_{n}(\hat{\theta})\|_{2}^{\alpha}+2\left(4\hat{R}_{n}(\hat{\Xi}^{(n)})+c_{3}\frac{\log\frac{1}{\delta}}{n}\right)^{\alpha}\bigg{)}+c_{4}\left(\frac{\log\frac{1}{\delta}}{n}\right)^{\frac{\alpha}{2}}$ (8)
For such a bound to be useful, the NERC $\hat{R}_{n}(\Xi^{(n)})$ needs to be small, e.g., $\hat{R}_{n}(\hat{\Xi}^{(n)})=O(\frac{1}{\sqrt{n}})$. [27] analyzed the NERC for loss functions of the form $\ell(\theta,(x,y))=h(\psi_{\theta}(x))$, where $\psi_{\theta}(x)=\langle\theta,x\rangle$ and $h$ is Lipschitz. For this case, it suffices to bound the NERC of the gradient of linear predictor $\psi_{\theta}(x)=\theta^{T}x$. Since $\nabla\psi_{\theta}(x)=x$ is independent of $\theta$, as long as each $\|x_{i}\|$ is bounded (Assumption 2), we have $\frac{1}{n}\mathbb{E}_{\varepsilon^{(n)}}\left[\sup_{\theta\in\Theta}\left\|\sum_{i=1}^{n}\varepsilon_{i}\nabla\phi(\theta,x_{i})\right\|\right]=\frac{1}{n}\mathbb{E}_{\varepsilon^{(n)}}\left[\left\|\sum_{i=1}^{n}\varepsilon_{i}x_{i}\right\|\right]=\frac{c}{\sqrt{n}}$, by Khintchine’s inequality [46, 82] for $L_{2}$ norm and under suitable assumptions on certain other norms [43, 42, 27]. The key simplification here is that the $\sup_{\theta\in\Theta}$ drops out since the gradient does not depend on $\theta$. However, the simplification is only possible for linear predictors. To go beyond, our analysis considers the general case where the $\sup_{\theta}$ does not drop out, and has to handled explicitly.

### v2

@@ Abstract
Generalization and optimization guarantees on the population loss often rely on uniform convergence based analysis, typically based on the Rademacher complexity of the predictors. The rich representation power of modern models has led to concerns about this approach. In this paper, we present generalization and optimization guarantees in terms of the complexity of the gradients, as measured by the Loss Gradient Gaussian Width (LGGW). First, we introduce generalization guarantees directly in terms of the LGGW under a flexible gradient domination condition, which includes the popular PL (Polyak-Łojasiewicz) condition as a special case. Second, we show that sample reuse in iterative gradient descent does not make the empirical gradients deviate from the population gradients as long as the LGGW is small. Third, focusing on deep networks, we bound their single-sample LGGW in terms of the Gaussian width of the featurizer, i.e., the output of the last-but-one layer. To our knowledge, our generalization and optimization guarantees in terms of LGGW are the first results of its kind, and hold considerable promise towards quantitatively tight bounds for deep models.
@@ 1 Introduction
Machine learning theory typically characterizes generalization behavior in terms of uniform convergence guarantees, often in terms of Rademacher complexities of predictor function classes [45, 10, 74, 61]. However, given the rich representation power of modern deep models, it has been difficult to develop sharp uniform bounds, and concerns have been raised whether non-vacuous bounds following the uniform convergence route may be possible [62, 63, 44, 43, 12, 32]. On the other hand, the empirical successes of deep learning has highlighted the importance of gradients, which play a critical role in learning over-parameterized models with deep representations [15, 33, 46]. The motivation behind this paper is to investigate generalization and optimization guarantees based on uniform convergence analysis of gradients, rather than the predictors.
@@
Empirical work in recent years has demonstrated that gradients of overparameterized deep learning models are typically “simple,” e.g., only a small fraction of entries have large values [65, 66, 48, 87, 89], gradients typically lie in a low-dimensional space [34, 30, 72, 40], using top-$k$ components with suitable bias corrections work well in practice [72, 40], etc. In this paper, we view such simplicity in terms of a generic chaining based “covering” of the gradients, which can be represented as the Loss Gradient Gaussian Width (LGGW) [81, 84]. Based on the empirical evidence, we posit that loss gradients of modern machine learning models have small LGGW. Our main technical results in the paper present generalization and optimization guarantees in terms of the LGGW. While our main technical results hold irrespective of whether the LGGW is small or not, similar to their classical Rademacher complexity counterparts, the results imply sharper bounds if the LGGW is small. Further, we consider deep learning models, specifically Feed-Forward Networks (FFNs) and Residual Networks (ResNets), and show that under standard assumptions, the LGGW of these models can be expressed in terms of the Gaussian width of the featurizer, i.e., the output of the last-but-one layer.
@@
In Section 2, we present generalization bounds in terms of the LGGW, based on a generic chaining argument over the gradients. Our analysis needs a mechanism to transition from losses to gradients of losses, and we assume the population loss function satisfies a flexible gradient domination (GD) condition [28], which includes the PL (Polyak-Łojasiewicz) condition [67] used in modern deep learning [51, 54, 52] as a simple special case. Then, based on recent advances in uniform convergence of vector-valued functions [57, 58, 28], one can readily get generalization bounds based on the vector Rademacher complexity of gradients. Our main new result on generalization is to show that such vector Rademacher complexity can be bounded in terms of the LGGW, which makes the bound explicitly depend on the geometry of gradients. Our proof is based on generic chaining (GC) [81], where we first bound the vector Rademacher complexity using a hierarchical covering argument, and then bound the hierarchical covering in terms of the LGGW using the majorizing measure theorem [81]. To our knowledge, this is the first generalization bound in terms of Gaussian width of loss gradients, and the results imply that models with small LGGW will generalize well.
@@
In Section 3, we present optimization guarantees based on the LGGW. One big disconnect between the theory and practice of (stochastic) gradient descent (GD) is that the theory assumes fresh samples in each step whereas in practice one reuses samples. If one just focuses on the finite sum empirical risk minimization (ERM), sample reuse is not an issue. However, the goal of such optimization is to reduce the population loss, and sample reuse derails the standard analysis. We show that even with sample reuse, the discrepancy between the sample average and the population gradient can be bounded by the LGGW along with only a logarithmic dependence on the number of steps of optimization, i.e., the number of times the samples has been reused, and also a logarithmic dependence on the ambient dimensionality. Thus, if the LGGW is small, the sample average gradients stay close to the population gradients, even under sample reuse. We also use these results to establish population convergence results for GD with sample reuse.
In Section 4, we present the first results on bounding LGGW of deep learning gradients. We consider formal models for FFNs and ResNets widely studied in the deep learning theory literature [1, 22, 8, 52, 9] and show that their single-sample LGGW can be bounded by the Gaussian width of the featurizer, i.e., the output of the last-but-one layer. We also demonstrate benefits of architecture choices, e.g., hidden layers being wider, last-but-one layer being narrower, which yields smaller single-sample LGGW.
@@
Notation: $c_{1},c_{2},c_{3}$, etc., denote constants, and they may mean different constants at different places. We use $\tilde{O}$ to hide poly-log terms.
@@ 2 Gaussian Width based Generalization Bounds
In this section, we formalize the notion of Loss Gradient Gaussian Width (LGGW) [84, 81], and as the main result, establish generalization bounds in terms of LGGW. For parameters $\theta\in\Theta\subset\mathbb{R}^{p}$ and data points $z\in{\cal Z}$ where $z=(\mathbf{x},y),\mathbf{x}\in\mathbb{R}^{d},y\in\mathbb{R}$, for any suitable loss $\ell(\theta;z):\Theta\times{\cal Z}\mapsto\mathbb{R}_{+}$, the gradients of interest are $\xi(\theta;z)=\nabla_{\theta}\ell(\theta;z)\in\mathbb{R}^{p}$. Given a dataset of $n$ samples, we start by defining the sets of all $n$-tuple of individual gradients and scaled stacked gradients:
@@
Given $n$ samples $z^{(n)}=\{z_{1},\ldots,z_{n}\}\in{\cal Z}^{(n)}$, for a suitable choice of parameter set $\Theta$ and loss $\ell:\Theta\times{\cal Z}\mapsto\mathbb{R}_{+}$, let the set of all possible $n$-tuple of individual loss gradients and scaled stacked gradients be respectively denoted as:
$\hat{\Xi}^{(n)}(\Theta)$ :=$:=\left\{\hat{\xi}^{(n)}=\{\xi_{i},i\in[n]\}\mid\exists\theta\in\Theta~{}\text{s.t.}~{}\xi_{i}=\nabla\ell(\theta;z_{i})\right\}~{},$ (1)
$\hat{\Xi}_{n}(\Theta)$ :=$:=\left\{\hat{\xi}_{n}\in\mathbb{R}^{np}\mid\exists\theta\in\Theta~{}~{}~{}\text{s.t.}~{}~{}\hat{\xi}_{n}=\frac{1}{\sqrt{n}}\big{[}\nabla\ell(\theta;z_{i})\big{]}_{i=1}^{n}\right\}~{},$ (2)
where $\big{[}\nabla\ell(\theta;z_{i})\big{]}_{i=1}^{n}$ denotes the $n$ $p$-dimensional gradients stacked to create $\hat{\xi}_{n}\in\mathbb{R}^{np}$.
The scaling of $\frac{1}{\sqrt{n}}$ in the stacked gradient vector $\hat{\xi}_{n}$ is to maintain the $L_{2}$-norm at the same scale as individual gradient components, i.e., with $\|\nabla\ell(\theta;z_{i})\|_{2}=O(1)$, we will have $\|\hat{\xi}_{n}\|_{2}=O(1)$.
For convenience, we will denote the set of $n$-tuples of gradients in (1) as $\hat{\Xi}^{(n)}$ and the scaled stacked gradients in (2) as $\hat{\Xi}_{n}$. Note that the sets $\hat{\Xi}^{(n)}$ and $\hat{\Xi}_{n}$ are conditioned on $z^{(n)}$, and the results we will present will hold for any $z^{(n)}\in{\cal Z}^{n}$, i.e., any $n$ samples under suitable regularity conditions (see Section 4.2.1). Next, we define the Loss Gradient Gaussian Width (LGGW):
@@
Given the scaled stacked gradient set $\hat{\Xi}_{n}$ as in (2) in Definition 1, the Loss Gradient Gaussian Width (LGGW) is defined as:
$w(\hat{\Xi}_{n}):=\mathbb{E}_{\mathbf{g}\sim{\cal N}(0,\mathbb{I}_{np\times np})}\left[\sup_{\hat{\xi}_{n}\in\hat{\Xi}_{n}}~{}\langle\hat{\xi}_{n},\mathbf{g}\rangle\right]~{}.$ (3)
We refer the reader to the book by Vershynin [84] for a technical introduction to Gaussian width and its properties. At a high level, Gaussian width measures how much elements of the set aligns with random Gaussian vectors, capturing the complexity of the set. As a special case of generic chaining (GC), Gaussian width estimates this alignment using expected supremum over Gaussian processes indexed on the set [81]. The GC perspective forms the basis of our technical development and we elaborate on this perspective further in the sequel.
@@ 2.1 Bounds based on Vector Rademacher Complexity
Our approach to generalization bounds is based on vector Rademacher complexities. Unlike the traditional approach which focus on bounds based on Rademacher complexities of the set of predictors $f(\theta;x)$ [45, 10, 74, 61], our work will leverage generalization bounds based on vector Rademacher complexities [57, 58, 28] of the set of loss gradients $\xi(\theta,z)=\nabla\ell(\theta;z)$. In particular, following [28], we consider Normed Empirical Rademacher Complexity (NERC), which is the extension of standard Rademacher complexity to sets of vectors or vector-valued functions.
@@
$\min_{\theta}{\cal L}_{{\cal D}}(\theta):=\mathbb{E}_{z\sim{\cal D}}\left[\ell(\theta;z)\right]~{}.$ (5)
The learner does not observe the distribution ${\cal D}$ and typically minimizes the empirical risk
^$\hat{\cal L}_{n}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\ell(\theta;z_{i})~{},$ (6)
@@
To understand whether the $(\alpha,\bar{c}_{\alpha})$-GD condition in Assumption 1 holds in deep learning models, in Figure 1, we plot the gradient domination ratio $\frac{{\cal L}_{\mathcal{D}}(\theta)}{\left\|\nabla{\cal L}_{\mathcal{D}}(\theta)\right\|_{2}^{\alpha}}$ for $\alpha=1,2$ for three standard deep learning models: ResNet, FFN, and CNN. The plots demonstrate that the condition indeed empirically holds in the optimization trajectory for $\alpha=1$ with a small constant $\bar{c}_{1}<5$ for all models. For $\alpha=2$, i.e., the PL condition [67, 51, 54, 52], the condition holds with $\bar{c}_{2}<1$ for some models, but needs large constants $\bar{c}_{2}>100$ for for ResNet18 on CIFAR-10. ∎
Remark 2.2.
The GD condition is more flexible than the PL condition as it allows for a more relaxed relationship between function values and the gradient norm. While both conditions may fail in landscapes with multiple valleys, starting optimization from a pre-trained model often positions the model within a single valley, where the GD condition is likely to hold. ∎
@@
Figure 1: Gradient Domination (GD) Ratio for $\alpha=1,2$ for (a) ResNet18 on CIFAR-10, (b) FFN on CIFAR-10, and (c) CNN on Fashion-MNIST. Each experiment is repeated 5 times and we report the average and the maximum ratio. The results show that GD holds for $\alpha=1$ with a small constant $\bar{c}_{1}<5$ for all models, and for $\alpha=2$, with $\bar{c}_{2}<1$ for some models, but needs large constants $\bar{c}_{2}>100$ for for ResNet18.
@@
${\cal L}_{{\cal D}}(\hat{\theta})-{\cal L}_{{\cal D}}(\theta^{*})\leq 2\bar{c}_{\alpha}\bigg{(}\|\nabla\hat{{\cal L}}_{n}(\hat{\theta})\|_{2}^{\alpha}+2\left(4\hat{R}_{n}(\hat{\Xi}^{(n)})+c_{3}\frac{\log\frac{1}{\delta}}{n}\right)^{\alpha}\bigg{)}+c_{4}\left(\frac{\log\frac{1}{\delta}}{n}\right)^{\frac{\alpha}{2}}~{}.$ (8)
For such a bound to be useful, the NERC $\hat{R}_{n}(\Xi^{(n)})$ needs to be small, e.g., $\hat{R}_{n}(\hat{\Xi}^{(n)})=O(\frac{1}{\sqrt{n}})$. [28] analyzed the NERC for loss functions of the form $\ell(\theta,(x,y))=h(\psi_{\theta}(x))$, where $\psi_{\theta}(x)=\langle\theta,x\rangle$ and $h$ is Lipschitz. For this case, it suffices to bound the NERC of the gradient of linear predictor $\psi_{\theta}(x)=\theta^{T}x$. Since $\nabla\psi_{\theta}(x)=x$ is independent of $\theta$, as long as each $\|x_{i}\|$ is bounded (Assumption 2), we have $\frac{1}{n}\mathbb{E}_{\varepsilon^{(n)}}\left[\sup_{\theta\in\Theta}\left\|\sum_{i=1}^{n}\varepsilon_{i}\nabla\phi(\theta,x_{i})\right\|\right]=\frac{1}{n}\mathbb{E}_{\varepsilon^{(n)}}\left[\left\|\sum_{i=1}^{n}\varepsilon_{i}x_{i}\right\|\right]=\frac{c}{\sqrt{n}}$, by Khintchine’s inequality [47, 83] for $L_{2}$ norm and under suitable assumptions on certain other norms [44, 43, 28]. The key simplification here is that the $\sup_{\theta\in\Theta}$ drops out since the gradient does not depend on $\theta$. However, simplification is only possible for linear predictors. Our analysis considers the general case where the $\sup_{\theta}$ does not drop out and has to be handled explicitly.